# Primitive immersions of constant curvature of surfaces into flag manifolds

### v1

@@ 3. Which invariant metric maximizes the area?
[…]
@@
[…]
is the degree [15, Chap.1] of the holomorphic map $\iota\circ\psi^{j}:S^{2}\to{\mathbb{C}}P^{{n\choose k^{(j)}}-1}$, being $\psi^{j}=\psi_{0}\oplus\ldots\oplus\psi_{j}$ (with rank $k^{(j)}$) and $\iota$ the Plücker embedding of $G_{k^{(j)}}({\mathbb{C}}^{n})$. Under the normalization $\lambda_{0}^{2}+\ldots+\lambda_{p-1}^{2}=1,$ we can now easily address the problem of finding the invariant metric on the flag manifold that maximizes the induced area of $S^{2}$. In view of (18), such invariant metric has parameters
@@
Let $\psi_{0}:S^{2}\to G_{2}({\mathbb{C}}^{5})$ be the holomorphic map locally spanned by
@@
Since $\delta_{0}=4$ and $\delta_{1}=2$, we see that, under the normalization $\lambda_{0}^{2}+\lambda_{1}^{2}=1$, the metric $\Psi^{*}g=\left(\lambda_{0}\gamma_{0}+\lambda_{1}\gamma_{1}\right)dzd\bar{z}$ which maximizes the area of $S^{2}$ is given by $\lambda_{0}=\frac{2}{\sqrt{5}}$, $\lambda_{1}=\frac{1}{\sqrt{5}}$. The curvature on $S^{2}$ induced by $\Psi$ from an invariant metric on $F_{2,2,1}$ is again given by (17). Figure 2 shows the graphs of the curvature $K_{\lambda_{0},\lambda_{1}}(\Psi)$ for different values of $\lambda_{0},\lambda_{1}$, as functions of the latitude angle $\varphi$ on $S^{2}$.
@@
(19) $\Psi^{*}ds^{2}_{F}=\sum_{j=0}^{n-1}k_{j}\gamma_{j}dzd\bar{z},\quad\mbox{where each $k_{j}$ is a positive integer.}$.
@@
for some constant $A\in\operatorname{U}(N)$, where $\hat{V}_{N}$ is the standard section of the Veronese map. For each positive integer $m$, set $[m]=\{0,1,\ldots,m-1\}$. Denote by $(\hat{\sigma}_{j})_{i}$ the $i$-th component of $\hat{\sigma}_{j}$, with $i\in[L_{j}]$. Comparing components of both sides of (20) we get, for each multiindex

### v2

@@ 4. Which invariant metric maximizes the area?
(18) $A(\Psi)=\frac{1}{2i}\sum_{j=0}^{p-1}\lambda_{j}\int_{S^{2}}\gamma_{j}d\bar{z}\wedge dz=\pi\sum_{j=0}^{p-1}\lambda_{j}\delta_{j},$
@@
$\delta_{j}=\frac{1}{2\pi i}\int_{S^{2}}\gamma_{j}d\bar{z}\wedge dz$
is the degree (see [6, §3.] and [15, Chap.1]) of the holomorphic map $\iota\circ\psi^{j}:S^{2}\to{\mathbb{C}}P^{{n\choose k^{(j)}}-1}$, being $\psi^{j}=\psi_{0}\oplus\ldots\oplus\psi_{j}$ (with rank $k^{(j)}$) and $\iota$ the Plücker embedding of $G_{k^{(j)}}({\mathbb{C}}^{n})$. Under the normalization $\lambda_{0}^{2}+\ldots+\lambda_{p-1}^{2}=1,$ we can now easily address the problem of finding the invariant metric on the flag manifold that maximizes the induced area of $S^{2}$. In view of (18), such invariant metric has parameters
@@
Let $\psi_{0}:S^{2}\to G_{2}({\mathbb{C}}^{5})$ be the constant curved holomorphic immersion locally spanned by
@@
Since $\delta_{0}=4$ and $\delta_{1}=2$, we see that, under the normalization $\lambda_{0}^{2}+\lambda_{1}^{2}=1$, the metric which maximizes the area of $S^{2}$ is given by $\Psi^{*}g=\left(\lambda_{0}\gamma_{0}+\lambda_{1}\gamma_{1}\right)dzd\bar{z}$ with $\lambda_{0}=\frac{2}{\sqrt{5}}$, $\lambda_{1}=\frac{1}{\sqrt{5}}$. The curvature on $S^{2}$ induced by $\Psi$ from an invariant metric on $F_{2,2,1}$ is again given by (17). Figure 2 shows the graphs of the curvature $K_{\lambda_{0},\lambda_{1}}(\Psi)$ for different values of $\lambda_{0},\lambda_{1}$, as functions of the latitude angle $\varphi$ on $S^{2}$.
@@
(19) $\Psi^{*}ds^{2}_{F}=\sum_{j=0}^{n-1}k_{j}\gamma_{j}dzd\bar{z},\quad\mbox{where each $k_{j}$ is a positive integer.}$
@@
for some constant $A\in\operatorname{U}(N)$, where $\hat{V}_{N}$ is the standard section of the Veronese map. For each positive integer $m$, set $m]=\{0,1,\ldots,m-1\}$. Denote by $(\hat{\sigma}_{j})_{i}$ the $i$-th component of $\hat{\sigma}_{j}$, with $i\in[L_{j}]$. Comparing components of both sides of (20) we get, for each multiindex
@@
Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.
Declarations
Conflict of interest
The authors declare no conflict of interest.